Find the cube root of 13824 by prime factorisation method.Find the cube root of 13824 by prime factorisation method.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

First let us prime factorise 13824:

13824 = 2 × 2 × 2 × 2 × 2 × 2 × 2 × 2 × 2 × 3 × 3 × 3

= 2^3× 2^3× 2^3× 3^3

3√13824 = 2 × 2 × 2 × 3 = 24


Thanks